In a study on cholesterol levels a sample of 12 men and women was chosen. The...

  1. In a study on cholesterol levels a sample of 12 men and women was chosen. The plasma cholesterol levels (mmol/L) of the subjects were as follows: 6.0, 6.4, 7.0, 5.8, 6.0, 5.8, 5.9, 6.7, 6.1, 6.5, 6.3, and 5.8. We assume that these 12 subjects constitute a simple random sample of a population of similar subjects. Find a 95% confidence interval for the population standard deviation of the plasma cholesterol levels.
Answer #1

36.00 ५०.96 Solukon Given that D--12 य 6.0 6.4 7.0 49.00 6.8 3364 6.0 36.00 6.8 33.64 59 34.81 6.7 44.89 6.0 37.2) 6.5 42.25Given that na 12. 8. 0.153561 95% confidence interval por 21-C 1-0.95 0.05. 0.975 1-3 1-0.0253 dif = n-1 = 12-1 x xa, dop 21.